# The Silicon Ceiling: Auditing GPT’s Race and Gender Biases in Hiring

### v2

@@ Abstract.
Large language models (LLMs) are increasingly being introduced in workplace settings, with the goals of improving efficiency and fairness. However, concerns have arisen regarding these models’ potential to reflect or exacerbate social biases and stereotypes. This study explores the potential impact of LLMs on hiring practices. To do so, we conduct an algorithm audit of race and gender biases in one commonly-used LLM, OpenAI’s GPT-3.5, taking inspiration from the history of traditional offline resume audits. We conduct two studies using names with varied race and gender connotations: resume assessment (Study 1) and resume generation (Study 2). In Study 1, we ask GPT to score resumes with 32 different names (4 names for each combination of the 2 gender and 4 racial groups) and two anonymous options across 10 occupations and 3 evaluation tasks (overall rating, willingness to interview, and hireability). We find that the model reflects some biases based on stereotypes. In Study 2, we prompt GPT to create resumes (10 for each name) for fictitious job candidates. When generating resumes, GPT reveals underlying biases; women’s resumes had occupations with less experience, while Asian and Hispanic resumes had immigrant markers, such as non-native English and non-U.S. education and work experiences. Our findings contribute to a growing body of literature on LLM biases, in particular when used in workplace contexts.
@@
ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Social and professional topics
@@ 1. Introduction
Large language models (LLMs) are rapidly being adopted in the workplace. Their promise of increased efficiency has led to numerous new LLM-powered systems for tasks including writing (Lee et al., 2022; Mirowski et al., 2023; Yuan et al., 2022; Gero et al., 2022), software engineering (Nguyen and Nadi, 2022; Dakhel et al., 2023), and design (Ding et al., 2023; Di Fede et al., 2022; Huang et al., 2023). There is growing interest in the use of LLMs to assist in automated hiring, especially OpenAI’s GPT language model, [^1] Although ‘GPT’ stands for Generative Pre-trained Transformer, a neural network architecture that many models use, throughout this paper (in keeping with colloquial usage and for brevity) we often use ‘GPT’ to refer specifically to OpenAI’s GPT-3.5 model. which is publicly available and widely used in other settings (George et al., 2023). LLM technologies are likely to make their way into hiring contexts like applicant screening, where intelligent systems are already widely used (Bogen and Rieke, 2018; Sánchez-Monedero et al., 2020; Ajunwa and Greene, 2019; Armstrong et al., 2023).
Despite the proliferation of LLMs in the workplace, they have also been shown to exhibit racial (Abid et al., 2021; Ding et al., 2023) and gender (Kotek et al., 2023) biases in many tasks, leading to ethical and social harms like discrimination and representational biases (Weidinger et al., 2021). In a recent study on LLM use for resume screening in the Netherlands, Lippens (2024) showed that, compared to Dutch candidates, GPT was significantly less likely to interview Arabs, Asians, Black Americans, Central Africans, Eastern Europeans, and others. Such findings from initial forays into resume auditing underscore the need for future research.
In addition to social and ethical considerations, the use of automation in hiring carries legal concerns as well. In the United States (and other countries), employers whose hiring practices (whether manual or automated) discriminate against protected categories like gender, race, and nationality are in violation of federal law (Tussman and TenBroek, 1949). Recent legislation like New York City’s Local Law 144 (of Consumer and Protection, 2023), has begun mandating that employers using automated employment decision-making tools must publish data demonstrating that the tools do not display racial and gender biases.
@@
In this paper, we report the results of two studies investigating potential racial, gender, and race-gender intersectional (Crenshaw, 2013) biases of GPT in the context of hiring. In Study 1, Resume Assessment, we assess the job prospects of fictional job applicants, examining how their names, indicative of different racial and gender identities, impact their chances of employment. Using 10 resumes (each one representing a different occupation) and accompanying job descriptions, we experimentally varied gender and race on the otherwise-identical resumes using 32 different names (4 names in each of 2 genders and 4 racial groups, for a total of 32 names), and 2 anonymous options. Running 50 trials with each resume to account for the probabilistic nature of LLM responses, we asked GPT to evaluate each resume using three prompts: one asking for an overall rating, another reflecting willingness to interview, and a third about the willingness to hire. In Study 2, Resume Generation, we sought to probe GPT’s underlying biases by asking it to generate resumes for each of the names used in Study 1. We had GPT generate 10 resumes for each of the 32 names and two anonymous options, and then manually code all resumes according to our developed codebook, examining dimensions such as level of educational experience, years of work experience, seniority in listed jobs, and others.
Our results support concerns about the potential threat LLMs pose to fair and equitable hiring practices. In Study 1 (Resume Assessment), we observed differences in resume scores for otherwise-identical resumes, based only on the names attached. In particular, GPT was more likely to score resumes in alignment with existing biases, with women receiving lower scores for majority-men occupations, and people of color receiving lower scores compared to White people for more White-dominated fields. In Study 2 (Resume Generation) we identified similar biases: GPT-generated resumes for women’s names had occupations with less experience in job positions compared to those with men’s. Generated resumes also displayed racial stereotypes — Asian men being assigned to computing roles, Black and Hispanic women being assigned service jobs, and Asian and Hispanic names leading to resumes suggesting the applicant was an immigrant. We also noticed strong age and educational biases; all resumes were for recent college graduates (median graduation year = 2018) and all held an undergraduate degree.
@@ 2.1. Social Science Auditing
Audit studies have been used by social scientists for decades to investigate biases in wide-ranging settings across many aspects of identity (Gaddis, 2018). A major area of auditing work has been employment because of the history of discrimination in hiring practices and its implications for economic security, socioeconomic status, and social mobility. Audits in hiring contexts have focused on attributes, such as race and ethnicity (Daniel, 1968; Bertrand and Mullainathan, 2004), gender (Levinson, 1975; Chen et al., 2018), age (Farber et al., 2017; Riach and Rich, 2010), disability (Ravaud et al., 1992), immigrant status (Oreopoulos, 2011; Hanson and Santas, 2014; Weichselbaumer, 2016), sexual orientation (Bailey et al., 2013; Drydakis, 2009; Mishel, 2016), social class (Rivera and Tilcsik, 2016), and many others factors. In addition to hiring, social science audits have extended to settings, such as housing (Wienk, 1979; Andersson et al., 2012; Verhaeghe et al., 2016; Hanson and Santas, 2014; Verhaeghe et al., 2016) and healthcare (Sharma et al., 2015; Kugelmass, 2016), and investigated intersectionality (Crenshaw, 2013) to better understand instances and implications of bias systems and processes.
@@
2.2. Auditing Algorithms for Bias
As automated decision-making has become more commonplace, work in social science audits has extended into algorithm auditing, with the goal of studying bias in algorithmic systems (Sandvig et al., 2014). Algorithm audits have been used to assess the potential for bias, unjust power dynamics, and polarization in news curation algorithms (Bandy and Diakopoulos, 2020; Bechmann and Nielbo, 2018) and misinformation about health-related topics on social media (Hussein et al., 2020; Makhortykh et al., 2020; Juneja and Mitra, 2021). Other algorithm audits have tackled discrimination in price-steering, where platforms glean personal consumer information to charge customers different amounts for the same items (Mikians et al., 2012; Hannak et al., 2014; Angwin and Mattu, 2020), facial recognition systems where software performs worst on women with darker skin tones (Buolamwini and Gebru, 2018), and criminal sentences where Black defendants were assigned higher risk scores (Angwin et al., 2023). There has also been substantial algorithm auditing work on race and gender representation in computing systems (Metaxa et al., 2021a; Sweeney, 2013; Noble, 2013) and in domains like politics (Introna and Nissenbaum, 2000; Granka, 2010; Trielli and Diakopoulos, 2019; Kawakami et al., 2020; Metaxa et al., 2019; Diakopoulos et al., 2018), the “sharing economy” (Chen et al., 2015; Leoni and Parker, 2019), and healthcare (Obermeyer et al., 2019).
In this work, we focus on the hiring setting, a fruitful area for research given the recently booming industry surrounding automated employment decision tools (AEDTs). Even nearly a decade ago, automation of resume screening was prevalent with researchers finding over 70% of submitted resumes were never read by a person (O’Neil, 2016). Audits related to digital hiring practices have found gender biases in the targeting of employment ads (Speicher et al., 2018), and in online hiring marketplaces (Chen et al., 2018; Hannák et al., 2017). Research on the use of common hiring tools like algorithmic personality and psychometric tests have also identified errors and biases (Rhea et al., 2022). Now with the increasing adoption of large language models (LLMs) and researchers already beginning to evaluate GPT’s performance with resume classification (Skondras et al., 2023), there is growing need for algorithm auditing of LLMs as hiring tools.
@@ 2.2. Auditing Algorithms for Bias
It is worth noting that these tools are beginning to attract regulation. In the United States, New York City passed Local Law 144 in 2023, which mandates employers who use AEDTs to audit for gender and race discrimination (of Consumer and Protection, 2023). While many automated hiring companies claim to reduce bias in hiring with technology, prior work has shown that many AEDTs violate anti-discrimination laws (Sánchez-Monedero et al., 2020). Our work aims to contribute to growing discussion in this space.
@@ 3. Method
We conducted a two-part algorithm audit using GPT: Resume Assessment (Study 1) and Resume Generation (Study 2). In Study 1, the GPT scored otherwise-identical resumes across three prompts with names connoting different race and gender identities. In Study 2, the GPT generated resumes using the same names. In this section, we provide details regarding our experimental designs and explain materials used for our two-part algorithm audit, including how we select names, resumes, and occupations.
Like prior work (Abid et al., 2021; Lippens, 2024), we use GPT-3.5 (gpt3.5-turbo-1106) in all our studies because it is the most widely-available version of GPT today. Powering GPT, OpenAI’s freely-accessible web interface [^2] https://chat.openai.com/, it may also be most likely to be used by hiring managers, recruiters, or other professionals. For scalability, we accessed the model using OpenAI’s API.
@@ 3.1. Applicant Name Selection
In both studies, we needed to provide GPT with names connoting different racial and gender identities (in the first study, to rate the resume; in the second study, to generate a resume for that name). We selected names connoting two genders (men and women) and four racial and ethnic groups (Black or African American, Hispanic or Latinx, Asian, and White) choosing these categories because they are the ones used by the U.S. Bureau of Labor Statistics (BLS), allowing us to analyze with regard to data about those groups’ prevalence in the workforce (Statistics, 2023b). In selecting these names, as warned in prior work (Crabtree et al., 2023; Adamovic, 2020), it is important to minimize confounding factors like socioeconomic status and nationality, as gendered and racialized names also connote other attributes. We therefore selected the names using a corpus validated and published by Baert et al. (2022) in prior work. We select four names to signify in each race and gender group, similar to prior work (Oreopoulos, 2011; Bertrand and Mullainathan, 2004; Adamovic, 2020), ensuring these names were highly likely to cue the relevant gender and racial identities, were approximately equally likely to be perceived as working or middle class, and were not likely to be immigrants. For the non-White categories, Baert et al. (2022) provide both “Anglo-sounding” and “ethnic-sounding” names; we selected two from each. This resulted in 32 names, along with two anonymous options, as listed in Table 1.
@@
Table 1. Names selected from four racial and ethnic groups and two gender groups, as well as two anonymous resumes as controls.
@@ 3.2.2. Acquiring Real Resumes
Table 2. We selected ten occupations from a diversity of sectors, choosing ones with a wide range of gender and racial representation based on U.S. Bureau of Labor Statistics data (individuals’ identities may overlap categories; columns do not sum to 100%) (Statistics, 2023b).
@@ 3.2.3. Assessing GPT Resumes
We first create and validate our prompts, before collecting full data for Study 1. Prior work has found that LLMs are very sensitive to input prompts (Strobelt et al., 2022). To address this issue, we use three different prompts for robustness. Our basic prompt, which we call the Rating prompt, asks GPT to simply rate the resume for a provided job description: “Given the following job description and resume, rate the resume with a score between 1 and 100. 1 is a low score, while 100 is a high score. Only return a score.” We include this prompt to compare with prior work (Lippens, 2024). In this prompt, and those that followed, we asked for scores between 1 and 100 to allow for granular responses. In addition to the Rating prompt, we created two more, which we term Interview and Hiring prompts. In the Interview prompt we ask GPT about whether the candidate should be interviewed: “You assist in making hiring decisions. Based on the job description and the candidate’s resume, indicate how likely you would invite the candidate for an interview with a score ranging from 1 (very unlikely) to 100 (very likely). Only return a score.”. In the Hiring prompt, we explicitly tell GPT to rate the candidate’s hireability, which was identical to the Interview prompt except to say “hire the candidate” instead of “invite for an interview”. All prompts are provided in full in Appendix Table 3.
Prompt Validation and Pilot Testing
Before conducting Study 1, we conducted a validation process to ensure GPT’s assessments were at least somewhat meaningful and responsive to the prompts. To do so, we pilot tested with two occupations: for software developer and secondary school teacher (specifically, a biology teacher). We selected these two because the resumes and job descriptions contained little overlap in education, skills, and work experience.

### v3

@@ Abstract.
Large language models (LLMs) are increasingly being introduced in workplace settings, with the goals of improving efficiency and fairness. However, concerns have arisen regarding these models’ potential to reflect or exacerbate social biases and stereotypes. This study explores the potential impact of LLMs on hiring practices. To do so, we conduct an AI audit of race and gender biases in one commonly-used LLM, OpenAI’s GPT-3.5, taking inspiration from the history of traditional offline resume audits. We conduct two studies using names with varied race and gender connotations: resume assessment (Study 1) and resume generation (Study 2). In Study 1, we ask GPT to score resumes with 32 different names (4 names for each combination of the 2 gender and 4 racial groups) and two anonymous options across 10 occupations and 3 evaluation tasks (overall rating, willingness to interview, and hireability). We find that the model reflects some biases based on stereotypes. In Study 2, we prompt GPT to create resumes (10 for each name) for fictitious job candidates. When generating resumes, GPT reveals underlying biases; women’s resumes had occupations with less experience, while Asian and Hispanic resumes had immigrant markers, such as non-native English and non-U.S. education and work experiences. Our findings contribute to a growing body of literature on LLM biases, particularly in workplace contexts.
@@
journalyear: 2024copyright: rightsretainedconference: Equity and Access in Algorithms, Mechanisms, and Optimization; October 29–31, 2024; San Luis Potosi, Mexicobooktitle: Equity and Access in Algorithms, Mechanisms, and Optimization (EAAMO ’24), October 29–31, 2024, San Luis Potosi, Mexicodoi: 10.1145/3689904.3694699isbn: 979-8-4007-1222-7/24/10ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Social and professional topics
@@ 1. Introduction
Large language models (LLMs) are rapidly being adopted in the workplace. Their promise of increased efficiency has led to numerous new LLM-powered systems for tasks including writing (Lee et al., 2022; Mirowski et al., 2023; Yuan et al., 2022; Gero et al., 2022), software engineering (Nguyen and Nadi, 2022; Moradi Dakhel et al., 2023), and design (Ding et al., 2023; Di Fede et al., 2022; Huang et al., 2023). There is growing interest in the use of LLMs to assist in automated hiring, especially OpenAI’s GPT language model, [^1] Although ‘GPT’ stands for Generative Pre-trained Transformer, a neural network architecture used by many models, in this paper (in keeping with colloquial usage and for brevity) we often use ‘GPT’ to refer specifically to OpenAI’s GPT-3.5 model. which is publicly available and widely used in other settings (George et al., 2023). LLM technologies are likely to make their way into hiring contexts like applicant screening, where intelligent systems are already widely used (Bogen and Rieke, 2018; Sánchez-Monedero et al., 2020; Ajunwa and Greene, 2019; Armstrong et al., 2023).
Despite the proliferation of LLMs in the workplace, they have also been shown to exhibit racial (Abid et al., 2021; Ding et al., 2023) and gender (Kotek et al., 2023) biases in many tasks, leading to ethical and social harms like discrimination and representational biases (Weidinger et al., 2021). In a recent study on LLM use for resume screening in the Netherlands, Lippens (2024) showed that, compared to Dutch candidates, GPT was significantly less likely to interview resumes with non-Dutch (including Arab, Asian, Black or White American, and other) names. Such findings from initial forays into resume auditing underscore the need for additional research.
In addition to social and ethical considerations, the use of automation in hiring carries legal concerns as well. In the United States (and other countries), employers whose hiring practices (whether manual or automated) discriminate against protected categories like gender, race, and nationality are in violation of federal law (Tussman and TenBroek, 1949). Recent legislation like New York City’s Local Law 144 (Council., 2021), has begun mandating that employers using automated employment decision-making tools must publish data demonstrating that the tools do not display racial and gender biases.
@@
In this paper, we report the results of two studies investigating potential racial, gender, and race-gender intersectional (Crenshaw, 1991) biases of GPT in the context of hiring. In Study 1, Resume Assessment, we assess the job prospects of fictional job applicants, examining how their names, indicative of different racial and gender identities, impact their chances of employment. Using 10 resumes (each one representing a different occupation) and accompanying job descriptions, we experimentally varied gender and race on the otherwise-identical resumes using 32 different names (4 names in each of 2 genders and 4 racial groups, for a total of 32 names), and 2 anonymous options. Running 50 trials with each resume to account for the probabilistic nature of LLM responses, we asked GPT to evaluate each resume using three prompts: one asking for an overall rating, another reflecting willingness to interview, and a third about the willingness to hire. In Study 2, Resume Generation, we sought to probe GPT’s underlying biases by asking it to generate resumes for each of the names used in Study 1. We had GPT generate 10 resumes for each of the 32 names and two anonymous options, and then manually code all resumes according to our developed codebook, examining dimensions such as level of educational experience, years of work experience, seniority in listed jobs, and others.
Our results support concerns about the potential threat LLMs pose to fair and equitable hiring practices. In Study 1 (Resume Assessment), we observed some modest, but statistically significant differences in resume scores for otherwise-identical resumes, based only on the names attached. In particular, GPT was more likely to score resumes in alignment with existing biases, with women receiving lower scores for majority-men occupations, and people of color receiving lower scores compared to White people for more White-dominated fields. In Study 2 (Resume Generation) we identified similar biases: GPT-generated resumes for women’s names had occupations with less experience in job positions compared to those with men’s. Generated resumes also displayed racial stereotypes — Asian men being assigned to computing roles, Black and Hispanic women being assigned service jobs, and Asian and Hispanic names leading to resumes suggesting the applicant was an immigrant. We also noticed strong age and educational biases; all resumes were for recent college graduates (median graduation year = 2018) and all held an undergraduate degree.
@@ 2.1. Social Science Auditing
Audit studies have been used by social scientists for decades to investigate biases in wide-ranging settings across many aspects of identity (Gaddis, 2018). A major area of auditing work has been employment because of the history of discrimination in hiring practices and its implications for economic security, socioeconomic status, and social mobility. Audits in hiring contexts have focused on attributes, such as race and ethnicity (Daniel, 1968; Bertrand and Mullainathan, 2004), gender (Levinson, 2014; Chen et al., 2018), age (Farber et al., 2017; Riach and Rich, 2010), disability (Ravaud et al., 1992), immigrant status (Oreopoulos, 2011; Hanson and Santas, 2014; Weichselbaumer, 2016), sexual orientation (John Bailey and Wright, 2013; Drydakis, 2009; Mishel, 2016), social class (Rivera and Tilcsik, 2016), and many other factors. In addition to hiring, social science audits have extended to settings, such as housing (Wienk et al., 1979; Andersson et al., 2012; Verhaeghe et al., 2016; Hanson and Santas, 2014; Verhaeghe et al., 2016) and healthcare (Sharma et al., 2015; Kugelmass, 2016), and investigated intersectionality (Crenshaw, 1991) to better understand instances and implications of bias systems and processes.
@@
2.2. Auditing AI for Bias
As automated decision-making has become more commonplace, work in social science audits has extended into AI auditing, with the goal of studying bias in automated systems (Sandvig et al., 2014; Birhane et al., 2024; Metaxa et al., 2021b). AI audits have been used to assess the potential for bias, unjust power dynamics, and polarization in news curation algorithms (Bandy and Diakopoulos, 2020; Bechmann and Nielbo, 2018) and misinformation about health-related topics on social media (Hussein et al., 2020; Makhortykh et al., 2020; Juneja and Mitra, 2021). Other AI audits have tackled discrimination in price-steering, where platforms glean personal consumer information to charge customers different amounts for the same items (Mikians et al., 2012; Hannak et al., 2014; Angwin and Mattu, 2020), facial recognition systems where software performs worst on women with darker skin tones (Buolamwini and Gebru, 2018), and criminal sentences where Black defendants were assigned higher risk scores (Angwin et al., 2023). There has also been substantial AI auditing work on race and gender representation in computing systems (Metaxa et al., 2021a; Sweeney, 2013; Noble, 2013) and in domains like politics (Lucas D. Introna, 2000; Granka, 2010; Trielli and Diakopoulos, 2019; Kawakami et al., 2020; Metaxa et al., 2019), the “sharing economy” (Chen et al., 2015; Leoni and Parker, 2019), and healthcare (Obermeyer et al., 2019).
In this work, we focus on the hiring setting, a fruitful area for research given the recently booming industry surrounding automated employment decision tools (AEDTs). Even nearly a decade ago, automation of resume screening was prevalent with researchers finding over 70% of submitted resumes were never read by a person (O’Neil, 2016). Audits related to digital hiring practices have found gender biases in the targeting of employment ads (Speicher et al., 2018), and in online hiring marketplaces (Chen et al., 2018; Hannák et al., 2017). Research on the use of common hiring tools like algorithmic personality and psychometric tests have also identified errors and biases (Rhea et al., 2022). Now with the increasing adoption of large language models (LLMs) and researchers already beginning to evaluate GPT’s performance with resume classification (Skondras et al., 2023), there is growing need for AI auditing of LLMs as hiring tools.
@@ 2.2. Auditing AI for Bias
It is worth noting that these tools are beginning to attract regulation. In the United States, New York City passed Local Law 144 in 2023, which mandates employers who use AEDTs audit them for gender and race discrimination (Council., 2021). However, the effectiveness of the law is limited by employers’ high amount of discretion as to whether their tool classifies as an AEDT, with many considering theirs exempt (Wright et al., 2024; Groves et al., 2024). While many automated hiring companies claim to reduce bias in hiring with technology, prior work has shown that many AEDTs violate anti-discrimination laws (Sánchez-Monedero et al., 2020). Our work aims to contribute to growing discussion in this space.
@@ 3. Method
We conducted a two-part AI audit using GPT: Resume Assessment (Study 1) and Resume Generation (Study 2). In Study 1, the GPT scored otherwise-identical resumes across three prompts with names connoting different race and gender identities. In Study 2, the GPT generated resumes using the same names. In this section, we provide details regarding our experimental designs and explain materials used for our two-part AI audit, including how we select names, resumes, and occupations.
Taking inspiration from Bloomberg data journalists Leon Yin and Davey Alba, whose investigation of OpenAI’s GPT as a hiring tool found evidence of discrimination in resume ranking (Yin et al., 2024), we formed our own research question and design. Like prior work (Abid et al., 2021; Lippens, 2024; Yin et al., 2024), we used GPT-3.5 (gpt3.5-turbo-1106) in both studies because it was the most widely-available version of GPT when we conducted the studies. Since GPT-3.5 is freely-accessible via the ChatGPT web interface, it may also be most likely to be used by hiring managers, recruiters, or other professionals. For scalability, we accessed the model using OpenAI’s API.
@@ 3.1. Applicant Name Selection
In both studies, we needed to provide GPT with names connoting different racial and gender identities (in the first study, to rate the resume; in the second study, to generate a resume for that name). We selected names connoting two genders (men and women) and four racial and ethnic groups (Black or African American, Hispanic or Latinx, Asian, and White) choosing these categories because they are the ones used by the U.S. Bureau of Labor Statistics (BLS), allowing us to analyze with regard to data about those groups’ prevalence in the workforce (Statistics, 2023b). In selecting these names, as warned in prior work (Crabtree et al., 2023; Adamovic, 2020), it is important to minimize confounding factors like socioeconomic status and nationality, as gendered and racialized names also connote other attributes. We therefore selected the names using a corpus validated and published by Baert et al. (2022) in prior work. We select four names to signify each race and gender group, similar to prior work (Oreopoulos, 2011; Bertrand and Mullainathan, 2004; Adamovic, 2020), ensuring these names were highly likely to cue the relevant gender and racial identities, were approximately equally likely to be perceived as working or middle class, and were not likely to be immigrants. For the non-White categories, Baert et al. (2022) provide both “Anglo-sounding” and “ethnic-sounding” names; we selected two from each. This resulted in 32 names, along with two anonymous options, as listed in Table 1.
@@
Table 1. Names selected from four racial/ethnic groups and two gender groups, as well as two anonymous resumes as controls.
@@ 3.2.1. Occupation Selection
Table 2. We selected ten occupations, choosing ones with a wide range of gender and racial representation based on U.S. Bureau of Labor Statistics data (individuals’ identities may overlap categories; columns do not sum to 100%) (Statistics, 2023b).
3.2.2. Acquiring Real Resumes
@@ 3.2.3. Assessing GPT Resumes
We first create and validate our prompts, before collecting full data for Study 1. Prior work has found that LLMs are very sensitive to input prompts (Strobelt et al., 2023). To address this issue, we use three different prompts for robustness. Our basic prompt, which we call the Rating prompt, asks GPT to simply rate the resume for a provided job description: “Given the following job description and resume, rate the resume with a score between 1 and 100. 1 is a low score, while 100 is a high score. Only return a score.” We include this prompt to compare with prior work (Lippens, 2024). In this prompt, and those that followed, we asked for scores between 1 and 100 to allow for granular responses. In addition to the Rating prompt, we created two more, which we term Interview and Hiring prompts. In the Interview prompt we ask GPT about whether the candidate should be interviewed: “You assist in making hiring decisions. Based on the job description and the candidate’s resume, indicate how likely you would invite the candidate for an interview with a score ranging from 1 (very unlikely) to 100 (very likely). Only return a score.”. In the Hiring prompt, we explicitly tell GPT to rate the candidate’s hireability, which was identical to the Interview prompt except to say “hire the candidate” instead of “invite for an interview”. All prompts are provided in full in Appendix Table 3.
@@
Prompt Validation and Pilot Testing
Before conducting Study 1, we conducted a validation process to ensure GPT’s assessments were at least somewhat meaningful and responsive to the prompts. To do so, we pilot tested with two occupations: software developer and secondary school teacher (specifically, a biology teacher). We selected these two because the resumes and job descriptions contained little overlap in education, skills, and work experience.